OpenStudy (anonymous):

-21x/7-5x=24 Solve the Equation. Pls explain.

OpenStudy (anonymous):

x = -3

OpenStudy (anonymous):

−21x 7 −5x=24 −3x+−5x=24 (−3x+−5x)=24(Combine Like Terms) −8x=24 −8x=24

OpenStudy (anonymous):

:( Pepe Pls

OpenStudy (anonymous):

Divide both sides by -8. −8x −8 = 24 −8 x=−3

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

I don't get the first part. What did you do to 21x/7

OpenStudy (anonymous):

well 21x/7 - 5x = 24

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

i shall await you here

OpenStudy (anonymous):

ok so let make 21 3 so -3x-5x=?

OpenStudy (anonymous):

still here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so i divide a coeficient with a constant. and i get - 3x - 5x = 24. Ok ty so much that makes it a lot easier. I did not know i could do that.

OpenStudy (anonymous):

X=-3

OpenStudy (anonymous):

TY gr8 teaching